OpenStudy (anonymous):

What are the roots of 3x^2 + 10 = 4x?

OpenStudy (anonymous):

mathway.com can help you with this

OpenStudy (anonymous):

Thanks Ash.("

OpenStudy (anonymous):

but make sure you type it in right atleast 2 times to check yourself to see if you get the same answer..

OpenStudy (trojanpoem):

a = 3 b = -4 c = 10\[\frac{ 4 \pm \sqrt{(-4)^2 - (4 * 3 * 10)} }{ 2 * 3 }\] \[\frac{ 4 \pm 2\sqrt{-26} }{ 6 } = \frac{ 2 \pm \sqrt{26}i }{ 3 }\]